A dehydrated patient needs a 8.15% saline IV. Unfortunately the hospital only has bags of 8% and 10% saline solutions. How many liters of each of these solutions should be mixed together to yield 4 liters of the desires concentration?
I don't know where to start.

Answer by stanbon(75887)

A dehydrated patient needs a 8.15% saline IV. Unfortunately the hospital only has bags of 8% and 10% saline solutions. How many liters of each of these solutions should be mixed together to yield 4 liters of the desires concentration?
----
Equation:
salt + salt = salt
0.08x + 0.10(4-x) = 0.0815*4
----
0.08x + 0.40 - 0.10x = 0.0815*4
---
Multiply thru by 10000 to get:
800x + 4000 - 1000x = 815*4
---
-200x = -740
x = 3.7 liters (amt. of 8% solution needed)
----
4-x = 4 - 3.7 = 0.3 liters (amt. of 10% solution needed)
=================================
Cheers,
Stan H.
==============